OpenStudy (anonymous):

solve the equation (t+5/8)-(t-2/2)=1/3

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (anonymous):

no clue

OpenStudy (skullpatrol):

$$\Huge \frac{t+5}{8}-\frac{t-2}{2}=\frac{1}3$$

OpenStudy (skullpatrol):

Is this the question?

OpenStudy (anonymous):

yes

OpenStudy (skullpatrol):

What is the common denominator of 2 and 8?

OpenStudy (anonymous):

2

OpenStudy (skullpatrol):

The LCM?

OpenStudy (skullpatrol):